OpenStudy (anonymous):

Circle S has an equation of (x – 16)2 + (y – 9)2 = 4. What is the center and radius of circle S? Answer Center: (16, 9); Radius: 4 Center: (-16, -9); Radius: 4 Center: (16, 9); Radius: 2 Center: (-16, -9); Radius: 2

OpenStudy (anonymous):

the third one.

OpenStudy (rulnick):

madsci is right

OpenStudy (anonymous):

can you explain? i really dont get it

OpenStudy (anonymous):

circle equation: (x-h)^2+(y-k)^2=r^2 h is the x coordinate, for the center, and k is the y coordinate for the center. r is for radius. Since (x-16)^2+(y-9)^2=4 we can say that r^2=4 so in order to find r we find the positive square root of four... which is two... so r=2 we know that: -16=-h multiply both sides by negative 1 and you get 16=h do the same for -9=-k and you get k=9 therefor, the coordinates for the center are (16,9) and the radius is 2

OpenStudy (anonymous):

oh ok i understand thanks so much

OpenStudy (anonymous):

no prob :)